# Disentangling Latent Shifts of In-Context Learning with Weak Supervision

### v2

@@ 1 Introduction
To address these limitations, adopting a mechanistic perspective on ICL has proven useful. In this view, demonstrations influence model behavior by inducing latent shifts – context-dependent changes in internal representations that alter how the model processes the query. By disentangling these shifts from the representation of the query itself, ICL can operate more robustly, processing queries independently of demonstrations. This, in turn, enables contextual knowledge to be stored persistently, eliminating the need to reprocess demonstrations for each new query. The latent shifts can then be reapplied directly, thereby reducing prompt length, improving inference efficiency, and enabling more modular and reusable representations. The disentanglement of demonstration-induced latent shifts from those of the query has been explored from both practical and theoretical perspectives: some work aims to improve the stability and scalability of ICL [27, 41], while other studies provide formal insights into the nature of context-induced shifts [7, 33]. However, existing methods rely on approximations by manipulating attention heads or hidden states. Developing a more direct approach to disentangling these shifts remains a compelling direction for future work.
In this paper, we propose to disentangle the latent shifts of the demonstrations and the query by using the output of ICL as a weak supervision signal, rather than approximating the underlying internal mechanisms. Specifically, we use ICL predictions as pseudo-labels that capture the full, contextualized influence of demonstrations. These pseudo-labels guide the training of a student model that internalizes demonstration-conditioned behavior – without repeated prompting or architectural intervention. We instantiate this idea with wilda (Weakly-supervised In-context Learning Disentanglement via Adapters), a parameter-efficient method for encoding the latent shifts induced by in-context demonstrations. In a teacher–student setup, an ICL-based teacher generates pseudo-labels for unlabeled queries, while the student learns to predict them using only the query input. The student updates a lightweight adapter module [15], enabling it to capture the shift in a reusable and modular form. Once trained, the adapter supports efficient inference without requiring demonstrations in the prompt. At inference time, additional demonstrations can still be provided, and are composed with the adapter’s shift – enabling flexible combinations of prompt-based and parameter-based task control.
@@
Our contribution is twofold: (1) we propose wilda, a method that encodes ICL-induced behavior into reusable adapters, improving inference efficiency and prompt stability without requiring demonstration prompts; and (2) we show that wilda outperforms traditional ICL and latent shift disentanglement methods on both ID and OOD data, while maintaining parameter efficiency. Together, these results show that treating ICL as a source of weak supervision enables robust and scalable behavior modeling – offering an effective alternative to ICL inference.
@@ 3.1 Experimental Setup
We assess model performance on seven tasks from the GLUE benchmark [36], covering single-sequence binary classification (cola, sst, rte), sequence-pair binary classification (mrpc, qqp, qnli), and sequence-pair multi-class classification (mnli). Evaluation metrics follow established standards: Matthew’s correlation for cola, $F_{1}$ scores for mrpc and qqp, and accuracy for the remaining tasks, with evaluations conducted on the development sets. Additionally, we measure accuracy on selected datasets from the MMLU benchmark [13], specifically “elementary math” (math) and “miscellaneous” (misc). Predictions are made based on the probability of generating specific verbalizer tokens as the first token output by the models, facilitated by carefully crafted prompts designed explicitly for single-token answers (see Appendix F for detailed templates).
@@ 3.2 Generalization and Stability
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) 0-shot $62.3$ $79.1$ $64.3$ $59.9$ $44.6$ $63.6$ $61.1$ $31.5$ $62.5$ $n$-shot $75.1_{6.5}$ $93.5_{2.0}$ $77.0_{5.5}$ $68.0_{3.0}$ $58.5_{4.0}$ $74.0_{2.5}$ $70.0_{3.0}$ $43.5_{3.5}$ $84.0_{4.0}$ PBFT $73.2_{3.8}$ $93.8_{1.5}$ $77.8_{6.0}$ $67.4_{3.5}$ $56.5_{3.0}$ $72.0_{2.0}$ $68.0_{2.5}$ $44.0_{3.8}$ $83.5_{4.5}$ ICV $72.9_{2.7}$ $92.2_{1.8}$ $74.5_{6.3}$ $67.0_{4.2}$ $57.3_{3.5}$ $73.4_{2.3}$ $69.1_{2.8}$ $41.5_{4.3}$ $67.0_{4.2}$ Batch-ICL $77.8_{4.7}$ $94.1_{2.2}$ $78.0_{6.0}$ $70.9_{3.5}$ $59.8_{3.7}$ $75.2_{2.2}$ $\underline{72.5}_{2.7}$ $36.2_{4.0}$ $81.0_{2.5}$ wilda-F $83.4_{0.3}$ $95.1_{\textbf{0.6}}$ $\underline{80.3}_{\textbf{1.4}}$ $72.1_{2.5}$ $\underline{63.7}_{\textbf{1.5}}$ $76.2_{1.8}$ $71.9_{1.9}$ $\underline{46.0}_{2.3}$ $\underline{86.0}_{2.3}$ wilda-S $\underline{86.0}_{\textbf{0.6}}$ $\textbf{96.1}_{1.2}$ $\textbf{81.4}_{2.2}$ $\underline{73.1}_{\textbf{2.0}}$ $\textbf{64.3}_{2.2}$ $\textbf{77.7}_{\textbf{1.5}}$ $\textbf{73.1}_{\textbf{1.8}}$ $\textbf{49.5}_{\textbf{2.0}}$ $\textbf{88.0}_{\textbf{2.2}}$ wilda-R $\textbf{86.5}_{3.0}$ $\underline{95.5}_{0.8}$ $79.0_{4.3}$ $\textbf{73.5}_{3.0}$ $62.5_{2.8}$ $\underline{76.5}_{1.9}$ $72.0_{2.2}$ $44.0_{2.7}$ $85.5_{3.3}$ Phi 3 (mini 4k) 0-shot $60.6$ $78.3$ $61.1$ $58.1$ $43.7$ $63.1$ $57.8$ $29.5$ $52.0$ $n$-shot $72.1_{5.2}$ $90.6_{2.1}$ $75.6_{3.2}$ $65.3_{3.1}$ $55.5_{4.1}$ $71.1_{2.6}$ $66.2_{3.7}$ $37.5_{3.6}$ $75.5_{4.1}$ PBFT $70.6_{4.3}$ $90.9_{1.9}$ $73.6_{3.4}$ $63.6_{3.6}$ $53.6_{3.1}$ $69.6_{2.3}$ $64.6_{2.6}$ $36.5_{4.1}$ $73.5_{4.6}$ ICV $71.5_{3.1}$ $89.1_{2.1}$ $74.3_{3.2}$ $64.1_{4.1}$ $54.1_{3.6}$ $70.8_{2.4}$ $65.4_{2.9}$ $36.0_{4.6}$ $74.0_{4.3}$ Batch-ICL $75.3_{4.2}$ $91.2_{2.6}$ $76.6_{3.1}$ $67.1_{3.6}$ $56.1_{4.1}$ $72.6_{2.6}$ $67.3_{2.8}$ $38.0_{3.9}$ $76.0_{4.1}$ wilda-F $\underline{80.4}_{\textbf{1.2}}$ $92.1_{\textbf{1.6}}$ $\underline{78.2}_{\textbf{1.3}}$ $\underline{69.7}_{2.4}$ $\underline{59.5}_{2.5}$ $73.5_{2.1}$ $\underline{68.6}_{2.2}$ $\underline{40.5}_{3.2}$ $\underline{77.5}_{3.6}$ wilda-S $\textbf{82.4}_{1.1}$ $\textbf{93.2}_{\textbf{1.6}}$ $79.2_{1.4}$ $\textbf{70.4}_{\textbf{1.1}}$ $\textbf{60.7}_{\textbf{2.3}}$ $\textbf{74.1}_{\textbf{1.4}}$ $\textbf{69.6}_{\textbf{1.9}}$ $\textbf{41.5}_{\textbf{2.3}}$ $\textbf{78.0}_{\textbf{3.3}}$ wilda-R $79.0_{1.9}$ $\underline{92.6}_{2.0}$ $\textbf{79.6}_{2.9}$ $68.6_{3.9}$ $58.6_{2.9}$ $\underline{73.6}_{2.0}$ $68.1_{2.3}$ $39.5_{3.6}$ $77.0_{3.7}$
@@
We first evaluate the generalization and stability of wilda on ID data. Table 1 reports the 16-shot ID generalization scores along with standard deviations. Across all datasets and models, wilda-S consistently achieves the best generalization scores, outperforming standard ICL, PBFT, and the disentanglement methods ICV and Batch-ICL (cf. Table 5 in the Appendix for results with Llama 2). Compared to standard ICL, wilda-S shows absolute improvements ranging from $2.6\%$ to $11.9\%$ for Llama 3 and $2.5\%$ to $10.3\%$ for Phi 3, where the differences in scores are statistically significant across all datasets. [^2] We assess the statistical significance using a two-tailed Wilcoxon signed-rank test ($p<0.05$), applying the Holm-Bonferroni method for family-wise error rate correction. Similar patterns hold for $n\in\{4,8,32\}$, where wilda-S also surpasses standard ICL (cf. Table 6 in the Appendix for other $n$-shot setups). Additionally, when a larger set $\mathcal{D}_{\text{unlab}}$ is used, there is a marginal improvement in scores, while stability improves even further (see Table 7 in the Appendix). Notably, the improvements in generalization with wilda-S, compared to standard ICL (the teacher model in wilda), provide strong evidence that the student model is exhibiting W2S generalization; we provide a more detailed analysis of this phenomenon in Section 4. While the wilda-F and wilda-R variants show similar generalization scores to wilda-S, they typically exhibit higher variance. This makes wilda-S the preferred choice due to its greater stability with respect to demonstration selection, as it consistently improves upon standard $n$-shot ICL across all datasets and models. This is supported by the statistically significant differences in standard deviations on all datasets for Llama 3 and on all but qnli for Phi 3.333We test for significance using a two-tailed Levene’s test ($p<0.05$) and apply the Holm-Bonferroni method to correct for family-wise error rate.
@@
Having looked at stability with respect to demonstration selection, we now turn to a more focused evaluation of stability with respect to demonstration ordering. Table 2 reports the standard deviations across 50 runs, where the same set of demonstrations is used, but their order is shuffled for each run. Designed to adapt to shuffled demonstrations, wilda-S shows the highest stability to demonstration ordering, as evidenced by the smallest standard deviation. The stability improvements with wilda-S over standard ICL are statistically significant across all datasets.3
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc LLama 3 (8B) $n$-shot $4.81$ $1.62$ $4.19$ $2.22$ $3.04$ $1.81$ $2.03$ $2.52$ $2.87$ PBFT $2.71$ $1.14$ $4.53$ $2.69$ $2.27$ $1.57$ $1.82$ $2.70$ $3.22$ ICV $2.09$ $1.23$ $4.08$ $2.81$ $1.95$ $1.61$ $2.03$ $1.96$ $3.18$ Batch ICL $3.04$ $1.47$ $2.89$ $2.24$ $2.53$ $\underline{1.42}$ $1.74$ $2.51$ $2.59$ wilda-F $\underline{1.32}$ $\underline{0.72}$ $\underline{1.53}$ $\underline{1.83}$ $\underline{1.76}$ $1.54$ $\underline{1.38}$ $\underline{1.89}$ $\underline{2.07}$ wilda-S 0.22 0.53 1.04 1.21 1.28 0.73 1.14 1.22 0.97 wilda-R $2.04$ $1.34$ $2.47$ $2.05$ $1.85$ $1.48$ $1.64$ $2.03$ $2.51$
@@ 3.3 Adapter Arithmetic
Table 4: ID generalization scores of knowledge fusion for Llama 3. The scores are averaged over $10$ runs with standard deviations shown as subscripts. The table compares the effectiveness of knowledge fusion from 22, 44, and $8$ subsets of $16$ demonstrations. The highest scores are in bold.

### v3

@@ 1 Introduction
To address these limitations, a mechanistic perspective on ICL has proven useful. In this view, demonstrations influence model behavior by inducing latent shifts – context-dependent changes in internal representations that alter how the model processes the query. Disentangling these shifts from the representation of the query itself allows ICL to operate more robustly, processing queries independently of demonstrations. This, in turn, enables contextual knowledge to be stored persistently, eliminating the need to reprocess demonstrations for each new query. The latent shifts can then be reapplied directly, thereby reducing prompt length, improving inference efficiency, and enabling more modular and reusable representations. The disentanglement of demonstration-induced latent shifts from those of the query has been explored from both practical and theoretical perspectives: some work aims to improve the stability and scalability of ICL [28, 42], while other studies provide formal insights into the nature of context-induced shifts [8, 34]. However, existing methods rely on approximations that intervene directly in attention heads or hidden states. In contrast, we adopt a functional perspective on ICL, capturing demonstration-induced shifts implicitly through model outputs rather than architectural manipulation. This enables disentanglement to emerge as a learned, self-aligned process rather than one achieved through explicit manipulation of internal states.
In this paper, we propose to disentangle the latent shifts of the demonstrations and the query by using the output of ICL as a weak supervision signal, rather than approximating the underlying internal mechanisms. Specifically, we use ICL predictions as pseudo-labels that capture the full, contextualized influence of demonstrations. These pseudo-labels guide the training of a student model that internalizes demonstration-conditioned behavior – without repeated prompting or architectural intervention. We instantiate this idea with wilda (Weakly-supervised In-context Learning Disentanglement via Adapters), a parameter-efficient method for encoding the latent shifts induced by in-context demonstrations. In a teacher–student setup, an ICL-based teacher generates pseudo-labels for unlabeled queries, while the student learns to predict them using only the query input. The student updates a lightweight adapter module [16], enabling it to capture the shift in a reusable and modular form. The adapter supports efficient inference without requiring demonstrations in the prompt, and multiple adapters can be combined through simple arithmetic to fuse knowledge across demonstration subsets. At inference time, additional in-context demonstrations can be composed with the adapter’s shift, enabling flexible integration of prompt-based and parameter-based adaptation.
@@
Our contribution is threefold: (1) we propose wilda, a method that encodes ICL-induced behavior into reusable adapters, improving inference efficiency and prompt stability without requiring demonstration prompts; (2) we show that wilda outperforms traditional ICL and latent shift disentanglement methods on both ID and OOD data, while maintaining parameter efficiency; and (3) we demonstrate that multiple adapters can be combined through simple arithmetic operations, enabling scalable composition of demonstration subsets and efficient handling of long-context scenarios. Together, these results demonstrate that viewing ICL as a form of weak supervision enables parametric generalization that can also flexibly incorporate contextual adaptation. wilda thus offers a scalable and composable framework for stable task adaptation in LLMs. [^1] Our code is available at https://github.com/josipjukic/wilda.
@@ 3.1 Experimental Setup
We assess model performance on seven tasks from the GLUE benchmark [37], covering single-sequence binary classification (cola, sst, rte), sequence-pair binary classification (mrpc, qqp, qnli), and sequence-pair multi-class classification (mnli). Evaluation metrics follow established standards: Matthew’s correlation for cola, $F_{1}$ scores for mrpc and qqp, and accuracy for the remaining tasks, with evaluations conducted on the development sets. Additionally, we measure accuracy on selected datasets from the MMLU benchmark [14], specifically “elementary math” (math) and “miscellaneous” (misc). We further extend our analysis to the ARC-Challenge benchmark [7] to assess reasoning and multi-hop generalization, which we show in Appendix D.
Predictions are made based on the probability of generating specific verbalizer tokens as the first token output by the models, facilitated by carefully crafted prompts designed explicitly for single-token answers (see Appendix F for detailed templates).
@@ 3.2 Generalization and Stability
GLUE MMLU Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) 0-shot $62.3$ $79.1$ $64.3$ $59.9$ $44.6$ $63.6$ $61.1$ $31.5$ $62.5$ $n$-shot $75.1_{6.5}$ $93.5_{2.0}$ $77.0_{5.5}$ $68.0_{3.0}$ $58.5_{4.0}$ $74.0_{2.5}$ $70.0_{3.0}$ $43.5_{3.5}$ $84.0_{4.0}$ PBFT $73.2_{3.8}$ $93.8_{1.5}$ $77.8_{6.0}$ $67.4_{3.5}$ $56.5_{3.0}$ $72.0_{2.0}$ $68.0_{2.5}$ $44.0_{3.8}$ $83.5_{4.5}$ ICV $72.9_{2.7}$ $92.2_{1.8}$ $74.5_{6.3}$ $67.0_{4.2}$ $57.3_{3.5}$ $73.4_{2.3}$ $69.1_{2.8}$ $41.5_{4.3}$ $67.0_{4.2}$ Batch-ICL $77.8_{4.7}$ $94.1_{2.2}$ $78.0_{6.0}$ $70.9_{3.5}$ $59.8_{3.7}$ $75.2_{2.2}$ $\underline{72.5}_{2.7}$ $36.2_{4.0}$ $81.0_{2.5}$ wilda-f $83.4_{0.3}$ $95.1_{\textbf{0.6}}$ $\underline{80.3}_{\textbf{1.4}}$ $72.1_{2.5}$ $\underline{63.7}_{\textbf{1.5}}$ $76.2_{1.8}$ $71.9_{1.9}$ $\underline{46.0}_{2.3}$ $\underline{86.0}_{2.3}$ wilda-s $\underline{86.0}_{\textbf{0.6}}$ $\textbf{96.1}_{1.2}$ $\textbf{81.4}_{2.2}$ $\underline{73.1}_{\textbf{2.0}}$ $\textbf{64.3}_{2.2}$ $\textbf{77.7}_{\textbf{1.5}}$ $\textbf{73.1}_{\textbf{1.8}}$ $\textbf{49.5}_{\textbf{2.0}}$ $\textbf{88.0}_{\textbf{2.2}}$ wilda-r $\textbf{86.5}_{3.0}$ $\underline{95.5}_{0.8}$ $79.0_{4.3}$ $\textbf{73.5}_{3.0}$ $62.5_{2.8}$ $\underline{76.5}_{1.9}$ $72.0_{2.2}$ $44.0_{2.7}$ $85.5_{3.3}$ Phi 3 (mini 4k) 0-shot $60.6$ $78.3$ $61.1$ $58.1$ $43.7$ $63.1$ $57.8$ $29.5$ $52.0$ $n$-shot $72.1_{5.2}$ $90.6_{2.1}$ $75.6_{3.2}$ $65.3_{3.1}$ $55.5_{4.1}$ $71.1_{2.6}$ $66.2_{3.7}$ $37.5_{3.6}$ $75.5_{4.1}$ PBFT $70.6_{4.3}$ $90.9_{1.9}$ $73.6_{3.4}$ $63.6_{3.6}$ $53.6_{3.1}$ $69.6_{2.3}$ $64.6_{2.6}$ $36.5_{4.1}$ $73.5_{4.6}$ ICV $71.5_{3.1}$ $89.1_{2.1}$ $74.3_{3.2}$ $64.1_{4.1}$ $54.1_{3.6}$ $70.8_{2.4}$ $65.4_{2.9}$ $36.0_{4.6}$ $74.0_{4.3}$ Batch-ICL $75.3_{4.2}$ $91.2_{2.6}$ $76.6_{3.1}$ $67.1_{3.6}$ $56.1_{4.1}$ $72.6_{2.6}$ $67.3_{2.8}$ $38.0_{3.9}$ $76.0_{4.1}$ wilda-f $\underline{80.4}_{\textbf{1.2}}$ $92.1_{\textbf{1.6}}$ $\underline{78.2}_{\textbf{1.3}}$ $\underline{69.7}_{2.4}$ $\underline{59.5}_{2.5}$ $73.5_{2.1}$ $\underline{68.6}_{2.2}$ $\underline{40.5}_{3.2}$ $\underline{77.5}_{3.6}$ wilda-s $\textbf{82.4}_{1.1}$ $\textbf{93.2}_{\textbf{1.6}}$ $79.2_{1.4}$ $\textbf{70.4}_{\textbf{1.1}}$ $\textbf{60.7}_{\textbf{2.3}}$ $\textbf{74.1}_{\textbf{1.4}}$ $\textbf{69.6}_{\textbf{1.9}}$ $\textbf{41.5}_{\textbf{2.3}}$ $\textbf{78.0}_{\textbf{3.3}}$ wilda-r $79.0_{1.9}$ $\underline{92.6}_{2.0}$ $\textbf{79.6}_{2.9}$ $68.6_{3.9}$ $58.6_{2.9}$ $\underline{73.6}_{2.0}$ $68.1_{2.3}$ $39.5_{3.6}$ $77.0_{3.7}$
@@
We first evaluate the generalization and stability of wilda on ID data. Table 1 reports the $16$-shot ID generalization scores along with standard deviations. Across all datasets and models, wilda-s consistently achieves the best generalization scores, outperforming standard ICL, PBFT, and the disentanglement methods ICV and Batch-ICL (cf. Table 6 in the Appendix for results with Llama 2). Compared to standard ICL, wilda-s shows absolute improvements ranging from $2.6\%$ to $11.9\%$ for Llama 3 and $2.5\%$ to $10.3\%$ for Phi 3, where the differences in scores are statistically significant across all datasets. [^3] We assess the statistical significance using a two-tailed Wilcoxon signed-rank test ($p<0.05$), applying the Holm-Bonferroni method for family-wise error rate correction. Similar patterns hold for $n\in\{4,8,32\}$, where wilda-s also surpasses standard ICL (cf. Table 8 in the Appendix for other $n$-shot setups). Additionally, when a larger set $\mathcal{D}_{\text{unlab}}$ is used, there is a marginal improvement in scores, while stability improves even further (see Table 9 in the Appendix). Notably, the improvements in generalization with wilda-s, compared to standard ICL (the teacher model in wilda), provide strong evidence that the student model is exhibiting W2S generalization; we provide a more detailed analysis of this phenomenon in Section 4. While the wilda-f and wilda-r variants show similar generalization scores to wilda-s, they typically exhibit higher variance. This makes wilda-s the preferred choice due to its greater stability with respect to demonstration selection, as it consistently improves upon standard $n$-shot ICL across all datasets and models. This is supported by the statistically significant differences in standard deviations on all datasets for Llama 3 and on all but qnli for Phi 3.444We test for significance using a two-tailed Levene’s test ($p<0.05$) and apply the Holm-Bonferroni method to correct for family-wise error rate.
@@
Having looked at stability with respect to demonstration selection, we now turn to a more focused evaluation of stability with respect to demonstration ordering. Table 2 reports the standard deviations across $50$ runs, where the same set of demonstrations is used, but their order is shuffled for each run. Designed to adapt to shuffled demonstrations, wilda-s shows the highest stability to demonstration ordering, as evidenced by the smallest standard deviation. The stability improvements with wilda-s over standard ICL are statistically significant across all datasets.4
@@
Beyond generalization and stability, we also assess whether the adapters faithfully encode the information contained in the demonstrations. To evaluate this, we encode single demonstrations into the adapters and measure the student model’s ability to reconstruct them using a simple recall task. As detailed in Appendix D.3, the adapters achieve consistently high semantic similarity across GLUE datasets, indicating that the demonstration content is reliably preserved. These findings corroborate the effectiveness of wilda in capturing and storing task-specific information within the adapter weights in a faithful and disentangled manner.
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc LLama 3 (8B) $n$-shot $4.81$ $1.62$ $4.19$ $2.22$ $3.04$ $1.81$ $2.03$ $2.52$ $2.87$ PBFT $2.71$ $1.14$ $4.53$ $2.69$ $2.27$ $1.57$ $1.82$ $2.70$ $3.22$ ICV $2.09$ $1.23$ $4.08$ $2.81$ $1.95$ $1.61$ $2.03$ $1.96$ $3.18$ Batch-ICL $3.04$ $1.47$ $2.89$ $2.24$ $2.53$ $\underline{1.42}$ $1.74$ $2.51$ $2.59$ wilda-f $\underline{1.32}$ $\underline{0.72}$ $\underline{1.53}$ $\underline{1.83}$ $\underline{1.76}$ $1.54$ $\underline{1.38}$ $\underline{1.89}$ $\underline{2.07}$ wilda-s 0.22 0.53 1.04 1.21 1.28 0.73 1.14 1.22 0.97 wilda-r $2.04$ $1.34$ $2.47$ $2.05$ $1.85$ $1.48$ $1.64$ $2.03$ $2.51$
@@ 3.3 Adapter Arithmetic
Table 4: ID generalization scores of knowledge fusion for Llama 3 (8B). The scores are averaged over $10$ runs with standard deviations shown as subscripts. The table compares the effectiveness of knowledge fusion from 22, 44, and $8$ subsets of $16$ demonstrations. The highest scores are in bold.
@@
3.4 Few-Shot wilda
The core idea behind wilda is to remove the need for explicit demonstrations by encoding their effect directly into the adapter. However, we also want to verify that standard ICL remains functional when the adapter is active. In particular, we examine whether the latent shift captured by the adapter composes additively with the contextual shift induced by new demonstrations provided during inference.
To this end, we evaluate wilda-s using Llama 3 (8B) in a mixed few-shot configuration. Specifically, we consider a setup in which the adapter has been fine-tuned to encode $16$ demonstrations and is further provided with an additional $16$ in-context demonstrations during inference. This configuration, denoted as $16/16$, thus combines parameter-based and context-based adaptation. We compare it against several baselines: standard $32$-shot ICL, wilda-s with $16$ encoded demonstrations and no in-context examples ($0/16$), and wilda-s with $32$ encoded demonstrations and no in-context examples ($0/32$).
The results, summarized in Table 5, show that the hybrid $16/16$ configuration outperforms standard $32$-shot ICL across all evaluated datasets, indicating that the adapter and in-context demonstrations reinforce one another rather than interfering. While $16/16$ setup performs slightly below the fully encoded $0/32$ variant, likely because the latter benefits from a dedicated fine-tuning phase, the $16/16$ setup demonstrates that wilda can successfully integrate additional ICL prompts on top of a fine-tuned adapter. This finding confirms that standard ICL remains fully compatible with adapter-based tuning, enabling a composition of contextual and parametric adaptation mechanisms within a single framework.
Table 5: Performance comparison of wilda-S configurations and standard 32-shot ICL using Llama 3 (8B). Results are averaged over 10 runs. The notation $n/d$ indicates the number of demonstrations in context ($n$) and encoded in the adapter ($d$).
GLUE MMLU Method rte sst qnli mnli cola mrpc qqp math misc $32$-shot ICL 75.3 93.2 77.7 69.1 58.3 76.4 74.2 43.0 84.5 wilda-s (0/32) 87.9 97.9 83.1 74.0 64.6 79.4 74.8 56.5 89.0 wilda-s (0/16) 86.0 96.1 81.4 73.1 64.3 77.7 73.1 49.5 88.0 wilda-s (16/16) 87.3 96.4 82.2 74.6 65.4 78.2 74.5 51.0 89.0